### v1

# Axial Bianchi I meets drifting extragalactic sources

Abstract
We recompute the drift of comoving sources in axial Bianchi I universes and correct the same calculation by Quercellini et al. from 2009. The correction is pertinent with respect to the Gaia catalogue of measured quasar drifts, that beg to be fitted.
Key-Words: cosmological parameters
1 Introduction
As early as 1933, Georges Lemaître [1], following a recommandation by Albert Einstein, considers (tri-axial) Bianchi I universes without using this name. They generalize flat Robertson-Walker universes by allowing three different scale factors in three orthogonal directions:
$\hbox{\rm{d}}\tau^{2}=\hbox{\rm{d}}t^{2}-a(t)^{2}\hbox{\rm{d}}x^{2}-b(t)^{2}\hbox{\rm{d}}y^{2}-c(t)^{2}\,\hbox{\rm{d}}z^{2},\quad a(t),\,b(t),\,c(t)>0$ (1)
These universes are homogeneous and anisotropic with a 3-dimensional Abelian isometry group. Its Lie algebra is the first in the list of Luigi Bianchi’s famous classification from 1898 [2] of the real, 3-dimensional Lie algebras. In the same work he also
  • exhibits 3-dimensional Riemannian manifolds on which these Lie algebras act transitively as Killing vectors,
  • he exhibits 3-dimensional Riemannian manifolds on which 4-dimensional Lie algebras act in the same way and shows that they all contain the 3-dimensional Lie algebras from his list as ideals,
  • and he shows that 5-dimensional Lie algebras cannot act in this way.
For Bianchi I this manifold is simply the 3-dimensional Eulidean space and the manifolds for the other Lie algebras of type Bianchi II - IX underlie the corresponding higher Bianchi universes, whose study began in the late sixties [3]-[5].
The universes with four Killing vectors are special cases, which are called locally rotationally symmetric (L.R.S.) or axial.
The proper time element of the axial Bianchi I universe is obtained by choosing two equal scale factors, e.g. $b=a$. Then there is a fourth Killing vector, $x\partial_{y}-y\partial_{x}$, generating rotations around the $z$-axis. If all three scale factors are equal, we are back at flat Robertson-Walker universes with the maximally symmetric Euclidean 3-space and three translations, three rotations.
The purpose of this work is an ab initio calculation of the drift of a comoving source as seen by a not necessarily comoving observer, whose peculiar velocity however is much smaller than the speed of light. We carefully separate the three parts of the calculation:
kinematics:
position of the source on the celestial sphere as a function of the observer’s cosmic time and the redshift of the source,
dynamics:
Einstein’s equation with cosmological constant and comoving dust,
approximation of small anisotropy:
$c(t)=a(t)[1+{\textstyle\frac{3}{2}}\eta(t)],\ \eta(t)\ll 1$.
Note that only axial Bianchi I and IX universes fulfill the three conditions: (i) next-to-maximal symmetry, (ii) compatibility with Einstein’s equation sourced by comoving dust and (iii) admitting maximal symmetry in the limit $\eta\rightarrow 0$.
In 2014 Darling [6] used tri-axial Bianchi I universes to fit the drift of 429 extra-galactic radio sources measured by Titov & Lambert [7] using Very Long Baseline Interferometry.
Today, thanks to the Gaia mission, we have hundreds of thousands of Quasars with drift and redshift measurements ($z$ up to 4.7) begging to be fitted.
2 The path of a light ray
The axial Bianchi I metric can be written
$\hbox{\rm{d}}\tau^{2}=\hbox{\rm{d}}t^{2}-a(t)^{2}\left[\hbox{\rm{d}}x^{2}+\hbox{\rm{d}}y^{2}\right]-c(t)^{2}\,\hbox{\rm{d}}z^{2},\quad a(t),\,c(t)>0.$ (2)
We follow notations and use results of [8]. Consider a geodesic $x^{\mu}(p)$ and denote the derivation with respect to the affine parameter $p$ by ˙:$\dot{}\vcentcolon=\hbox{\rm{d}}/\hbox{\rm{d}}p$. The four Killing vectors $\partial x,\,\partial y,\,\partial z$ and $x\partial y-y\partial x$ yield four conserved quantities by Emmy Noether’s theorem:
$A\vcentcolon=a(t(p))^{2}\dot{x},\ B\vcentcolon=a^{2}\dot{y},\ C\vcentcolon=c^{2}\dot{z}\quad\text{and}\quad Ay-Bx=0.$ (3)
The last equation implies that if $A$ and $B$ both vanish the geodesic is a straight line in the ±$\pm\,z$-direction. Otherwise the geodesic remains in the plane defined by the $z$-direction and the direction of the initial 3-velocity.
Let us suppose the geodesic is light-like. Then from $\dot{x}^{\mu}g_{\mu\nu}\dot{x}^{\nu}=0$ we obtain
$\dot{t}=\,\frac{1}{W}\,\quad\text{with}\quad W\vcentcolon=\left(\frac{A^{2}+B^{2}}{a(t)^{2}}\,+\,\frac{C^{2}}{c(t)^{2}}\right)^{-1/2}.$ (4)
Eliminating the affine parameter $p$ in favour of the cosmic time $t$ and using :${}^{\prime}\vcentcolon=\hbox{\rm{d}}/\hbox{\rm{d}}t$, the light-like geodesic must satisfy the three first order equations:
$\vec{x}^{\prime}=\,\frac{\dot{\vec{x}}}{\dot{t}}\,=\,W\begin{pmatrix}\,\frac{A}{a^{2}}\,\\[1.42262pt] \,\frac{B}{a^{2}}\,\\[1.42262pt] \,\frac{C}{c^{2}}\,\end{pmatrix}.$ (5)
Let us fix the Noether constants $A,\,B$ and $C$ in terms of the initial velocity of a photon emitted by a comoving source at time $t_{e}$ and position $\vec{x}_{e}\vcentcolon=\vec{x}(t_{e})$. To simplify notations we put $\vec{x}_{e}=0$.
Let us say that the photon arrives today $t=t_{0}$ in our telescope at position $\vec{x}_{0}\vcentcolon=\vec{x}(t_{0})$. Then we have:
$x_{0}=\int_{t_{e}}^{t_{0}}\,\frac{A}{a^{2}}\,W,\quad y_{0}=\int_{t_{e}}^{t_{0}}\,\frac{B}{a^{2}}\,W,\quad z_{0}=\int_{t_{e}}^{t_{0}}\,\frac{C}{c^{2}}\,W.$ (6)
In our conventions the speed of light is unity; proper time $\tau$, the coordinate time $t$ and the affine parameter $p$ have units of seconds, the comoving coordinates $x,\,y,\,z$ are dimensionless and the scale factors $a$ and $c$ and the Noether constants $A,\,B$ and $C$ carry seconds. Then $W$ is dimensionless. Equation (5) remains unchanged under a simultaneous rescaling of the three Noether constants. Therefore we may choose $N\vcentcolon=\sqrt{A^{2}+B^{2}+C^{2}}$ to be any positive constant. Through a rescaling of the coordinates $x,\,y,\,z$ we may achieve $a_{0}=c_{0},$ with $a_{0}\vcentcolon=a(t_{0}),\dots$
We also learn from equations (5) or (6) that in the Bianchi I metric, the apparent position of comoving sources in the sky changes with time, ‘cosmic parallax, apparent motion, proper motion or drift’ (unless they lie in the $xy$-plane, $C=0$, or on the $z$-axis, $A=B=0$).
3 A second, infinitesimally close light ray
Consider now a second photon emitted a short time $T$ later from the same source and arriving in our telescope a time $T_{D}$ after the first. (The subscript $\cdot_{D}$ stands for Doppler.) The following calculations are relevant in two physical situations:
  • $T$ is an atomic period at emission. Then $T_{D}$ is the Doppler shifted atomic period observed here and today and we define the redshift today
    $\underline{z}_{0}\vcentcolon=\,\frac{T_{D}-T}{T}\,.$ (7)
    We underline the redshift $\underline{z}$ to avoid confusion with the coordinate $z$.
    If we pretend to know the metric and it has nice properties, then the redshift allows us to compute the time of flight of the photon.
  • $T_{D}$ is a few years and we observe a change in the direction of the arriving photons. Of course, a change in direction can only be measured with respect to another direction that supposedly does not change. Axial Bianchi I universes have indeed the luxury of such fixed directions: the $z$-axis and the $xy$-plane.
Let us denote the second geodesic by $\tilde{\vec{x}}(t)$ with $t$ varying from $t_{\tilde{e}}\vcentcolon=t_{e}+T$ to $t_{\tilde{0}}\vcentcolon=t_{0}+T_{D}$.
In both situations the second photon is sent out at
$\tilde{\vec{x}}(t_{\tilde{e}})={\vec{x}}(t_{e})=\vcentcolon\vec{x}_{e}=0,$ (8)
because we suppose the source to be comoving. For the observer, on the other hand, we allow for a small peculiar velocity today
$\vec{V}=\vcentcolon a_{0}\,\vec{v}\,,\quad\quad V\ll 1.$ (9)
Then the photon arrives in her telescope at position $\vec{x}_{0}+T_{D}\,\vec{v}$:
$\tilde{\vec{x}}(t_{\tilde{0}})=\vcentcolon\tilde{\vec{x}}_{\tilde{0}}={\vec{x}}(t_{0})+T_{D}\,\vec{v}=\vec{x}_{0}+T_{D}\,\vec{v}.$ (10)
We suppose that $T$ is small with respect to the ‘time of flight’, $T\ll t_{0}-t_{e}$; so small that we may keep only linear terms in $T$ and $T_{D}$. Also the peculiar velocity is small enough so that we may keep only linear terms in $V$. In particular, we neglect the dilation of the observer’s proper time with respect to cosmic time $t$.
We may use a rotation in the $xy$-plane to bring the peculiar velocity into the form
$\vec{v}=\begin{pmatrix}v_{x}\\ 0\\ v_{z}\end{pmatrix}.$ (11)
Our task is to slightly change the initial direction of the second geodesic $\tilde{\vec{x}}(t)$ coded in the Noether constants $\tilde{A},\,\tilde{B}$ and $\tilde{C}$, in such a manner that $\tilde{\vec{x}}_{\tilde{0}}=\vec{x}_{0}+T_{D}\,\vec{v}$. To this end we set:
$\tilde{A}=A+\alpha N,\quad\tilde{B}=B+\beta N,\quad\tilde{C}=C+\gamma N,\quad\alpha,\,\beta,\,\gamma\sim T/(t_{0}-t_{e}).$ (12)
We may choose $\tilde{N}=N$ implying
$A\alpha+B\beta+C\gamma\sim 0.$ (13)
The second trajectory is computed as the first:
$\tilde{x}_{\tilde{0}}=\int_{t_{e}+T}^{t_{0}+T_{D}}\,\frac{\tilde{A}}{a^{2}}\,\tilde{W},\quad\tilde{y}_{\tilde{0}}=\int_{t_{e}+T}^{t_{0}+T_{D}}\,\frac{\tilde{B}}{a^{2}}\,\tilde{W},\quad\tilde{z}_{\tilde{0}}=\int_{t_{e}+T}^{t_{0}+T_{D}}\,\frac{\tilde{C}}{c^{2}}\,\tilde{W},$ (14)
with
[…]
We linearize $\tilde{W}$,
$\tilde{W}\!\!\!$ $\sim$ $\!\!\!W\left[1-W^{2}N\left(\frac{A\alpha+B\beta}{a^{2}}\,+\,\frac{C\gamma}{c^{2}}\right)\right]$ (16)
$\sim$ $W\left[1+W^{2}NC\gamma\left(\frac{1}{a^{2}}\,-\,\frac{1}{c^{2}}\right)\right]\sim W\left[1+\,\frac{N}{C}\,\gamma\left(N^{2}\,\frac{W^{2}}{a^{2}}\,-1\right)\right]$ (17)
$\sim$ $W\!\left[1-W^{2}N(A\alpha+B\beta)\!\!\left(\frac{1}{a^{2}}\,-\,\frac{1}{c^{2}}\right)\!\right]\sim\!W\left[1-\,\frac{NC}{A^{2}+B^{2}}\,\gamma\!\left(N^{2}\,\frac{W^{2}}{c^{2}}\,-1\right)\!\right],$ (18)
where we have used successively the two identities:
$W^{2}\left(\frac{1}{a^{2}}\,-\,\frac{1}{c^{2}}\right)=\,\frac{N^{2}}{C^{2}}\,\frac{W^{2}}{a^{2}}\,-\,\frac{1}{C^{2}}\,=-\,\frac{N^{2}}{A^{2}+B^{2}}\,\frac{W^{2}}{c^{2}}\,+\,\frac{1}{A^{2}+B^{2}}\,.$ (19)
Then we have to first order in $T/(t_{0}-t_{e})$:
$\tilde{x}_{\tilde{0}}\!\!\!$ $\sim$ $\!\!\!\int_{t_{e}+T}^{t_{0}+T_{D}}\left(\frac{\tilde{A}\tilde{W}}{a^{2}}\,+\,\frac{NW}{a^{2}}\,\alpha\right)$ (20)
$\sim$ $\!\!\!\,\frac{AW_{0}}{a^{2}_{0}}\,T_{D}-\,\frac{AW_{e}}{a^{2}_{e}}\,T+\!\!\int_{t_{e}}^{t_{0}}\left\{\frac{AW}{a^{2}}\,-\frac{AW}{a^{2}}\,\,\frac{NC}{A^{2}+B^{2}}\,\gamma\!\left(\frac{N^{2}W^{2}}{c^{2}}\,-1\right)+\,\frac{NW}{a^{2}}\,\alpha\right\}$ (21)
$\sim$ $\!\!\!x_{0}+A\,F\,T_{D}+\left(\frac{AC}{A^{2}+B^{2}}\,\gamma+\alpha\right)I-\,\frac{AC}{A^{2}+B^{2}}\,\gamma J.$ (22)
In (21) we have used the second member of (18) and in (22) we have introduced the short hands,
$F\vcentcolon=\,\frac{W_{0}}{a_{0}^{2}}\,-\,\frac{W_{e}}{a_{e}^{2}}\,\frac{T}{T_{D}}\,,\quad F_{c}\vcentcolon=\,\frac{W_{0}}{c_{0}^{2}}\,-\,\frac{W_{e}}{c_{e}^{2}}\,\frac{T}{T_{D}}\,,\quad I\vcentcolon=N\int_{t_{e}}^{t_{0}}\,\frac{W}{a^{2}}\,,\quad J\vcentcolon=N^{3}\int_{t_{e}}^{t_{0}}\,\frac{W^{3}}{a^{2}c^{2}}\,.$ (23)
Since $\tilde{x}_{\tilde{0}}=x_{0}+v_{x}T_{D}$ we have:
$v_{x}T_{D}$ $\sim$ $A\,F\,T_{D}+\alpha I+\,\frac{AC}{A^{2}+B^{2}}\,\gamma(I-J).$ (24)
Similarily for $\tilde{y}_{\tilde{0}}=y_{0}$ and $\tilde{z}_{\tilde{0}}=x_{0}+v_{z}T_{D}$:
00 0 $\sim$ $B\,F\,T_{D}+\beta I+\,\frac{BC}{A^{2}+B^{2}}\,\gamma(I-J),$ (25)
$v_{z}T_{D}$ $\sim$ $C\,F_{c}\,T_{D}+\gamma J.$ (26)
To derive equation (26) we used the second member of (17). Let us take the following linear combination of the last three equations, $A$(24)+$B$(25)+$C$(26):
$(Av_{x}+Cv_{z})\,T_{D}\sim(A^{2}+B^{2})\,F\,T_{D}+C^{2}\,F_{c}\,T_{D}=\,\frac{T_{D}}{W_{0}}\,-\,\frac{T}{W_{e}}\,.$ (27)
From it we obtain the direction-dependent redshift today,
$\underline{z}_{0}+1\vcentcolon=\,\frac{T_{D}}{T}\,\sim\,\frac{1}{W_{e}}\,\frac{1}{N/a_{0}-(Av_{x}+Cv_{z})}\,.$ (28)
We will also need the corrections $\alpha,\,\beta,\,\gamma$ to the second photon’s initial direction. To this end let us consider the linear combination $B$(24)$-A$(25):
$\alpha\sim\,\frac{A}{B}\,\beta+\,\frac{1}{I}\,v_{x}\,T_{D}.$ (29)
Using equations (13) and (26) we obtain,
$\alpha$ $\sim$ $\frac{1}{A^{2}+B^{2}}\,\frac{1}{J}\left[AC^{2}F_{c}+B^{2}\,\frac{J}{I}\,v_{x}-AC\,v_{z}\right]T_{D},$ (30)
$\beta$ $\sim$ $\frac{B}{A^{2}+B^{2}}\,\frac{1}{J}\left[C^{2}F_{c}+A\,\frac{J}{I}\,v_{x}-C\,v_{z}\right]T_{D},$ (31)
$\gamma$ $\sim$ $\hskip 48.36967pt\frac{1}{J}\left[-C\,F_{c}+v_{z}\right]T_{D}.$ (32)
4 Arriving angles
Our first task is to compute the angle $\theta$ between the $z$-axis and the velocity $\vec{x}^{\prime}_{0}$ of the first photon. To this end we use the negative of the spatial part of the metric (2). Note that the velocity vector $\vec{x}^{\prime}$ is a unit vector with respect to this 3-metric, because we have set the speed of light to one. If we denote by
$\vec{u}_{z}\vcentcolon=\begin{pmatrix}0\\ 0\\ \frac{1}{c}\end{pmatrix}$ (33)
the unit vector in the $z$-direction, we have
$\cos\theta=\vec{x}^{\prime}_{0}\cdot\vec{u}_{z0}=W_{0}\,\frac{C}{c_{0}^{2}}\,\,c_{0}^{2}\,\,\frac{1}{c_{0}}\,=\,\frac{CW_{0}}{c_{0}}\,=\,\frac{C}{N}\,.$ (34)
Our second task is to compute the same angle $\tilde{\theta}$ for the second photon at its arrival time $t_{\tilde{0}}=t_{0}+T_{D}$:
$\cos\tilde{\theta}={\tilde{\vec{x}}}^{\prime}_{\tilde{0}}\cdot\vec{u}_{z\tilde{0}}=\,\frac{\tilde{C}\tilde{W}_{\tilde{0}}}{c_{\tilde{0}}}\,.$ (35)
From the second expression (18) we obtain,
$\,\frac{\tilde{C}\tilde{W}}{c}\,\sim\,\frac{CW}{c}\left[1+\,\frac{N^{2}}{A^{2}+B^{2}}\,\frac{N}{C}\left(1-\left\{\frac{CW}{c}\right\}^{2}\right)\right],$ (36)
and
$\,\frac{\tilde{C}\tilde{W}_{\tilde{0}}}{c_{\tilde{0}}}\,\sim\,\frac{CW_{\tilde{0}}}{c_{\tilde{0}}}\left[1+\,\frac{N^{2}}{A^{2}+B^{2}}\,\frac{N}{C}\left(1-\left\{\frac{CW_{0}}{c_{0}}\right\}^{2}\gamma\right)\right]\sim\,\frac{CW_{\tilde{0}}}{c_{\tilde{0}}}\left[1+\,\frac{N}{C}\,\gamma\right].$ (37)
We Taylor develop the leading term $w(t)\vcentcolon=C\,W(t)/c(t)$ around $t_{0}$ viz
$w(t_{0}+T_{D})\sim w_{0}+T_{D}\,w^{\prime}_{0}$:
$w$ $=$ $\left(\frac{A^{2}+B^{2}}{C^{2}}\,\frac{c^{2}}{a^{2}}+1\right)^{-1/2}\,,$ (38)
$w^{\prime}$ $=$ $\,\frac{A^{2}+B^{2}}{C^{2}}\,\frac{c^{2}}{a^{2}}\,(H-H_{c})\left(\frac{A^{2}+B^{2}}{C^{2}}\,\frac{c^{2}}{a^{2}}+1\right)^{-3/2},\quad H\vcentcolon=\,\frac{a^{\prime}}{a}\,\quad H_{c}\vcentcolon=\,\frac{c^{\prime}}{c}\,,$ (39)
$w_{0}$ $=$ $\cos\theta,\quad\quad w_{0}^{\prime}\ =\ \sin^{2}\theta\,\cos\theta\,(H_{0}-H_{c0}).$ (40)
Therefore
$W_{\tilde{0}}\,\frac{C}{c_{\tilde{0}}}\sim\cos\theta\,[1+\sin^{2}\theta\,(H_{0}-H_{c0})\,T_{D}],$ (41)
and
$\cos\tilde{\theta}\sim\cos\theta\left[1+\sin^{2}\theta\,(H_{0}-H_{c0})\,T_{D}+\,\frac{\gamma}{\cos\theta}\right].$ (42)
The apparent angular drift of the comoving source $\delta\theta\vcentcolon=\tilde{\theta}-\theta$ then takes the form
$\delta\theta\sim-\sin\theta\,\cos\theta\,(H_{0}-H_{c0})\,T_{D}-\,\frac{\gamma}{\sin\theta}\,.$ (43)
We need a second angle to localize the direction of the incoming photons on the celestial sphere. Following standard conventions we choose the angle $\varphi$ in the $xy$-plane between the orthogonal projection of $\vec{x}^{\prime}$ into this plane and the positive $x$-axis. Denote this projection by $\vec{x}^{\prime}_{\perp}$ and the unit vector in the positive $x$-direction by $\vec{u}_{x}$,
$\vec{x}^{\prime}_{\perp}\vcentcolon=\,W\begin{pmatrix}\,\frac{A}{a^{2}}\,\\[1.42262pt] \,\frac{B}{a^{2}}\,\\[1.42262pt] 0\end{pmatrix},\quad\vec{u}_{x}\vcentcolon=\begin{pmatrix}\frac{1}{a}\\ 0\\ 0\end{pmatrix}.$ (44)
Then the angle $\varphi$ today is given by
$\cos\varphi=\,\frac{\vec{x}^{\prime}_{\perp 0}\cdot\vec{u}_{x0}}{\sqrt{\vec{x}^{\prime}_{\perp 0}\cdot\vec{x}^{\prime}_{\perp 0}}}\,=\,\frac{A}{\sqrt{A^{2}+B^{2}}}\,,$ (45)
and we retrieve the usual relations of spherical coordinates:
$A$ $=$ $N\cos\varphi\,\sin\theta,$ (46)
$B$ $=$ $N\sin\varphi\,\sin\theta,$ (47)
$C$ $=$ $N\quad\quad\,\cos\theta.$ (48)
Note that if $A=B=0$ the first photon moves along the $z$-axis, $\sin\theta=0$. Then the projection of the photon’s velocity onto the $xy$-plane vanishes and the angle $\varphi$ is not defined.
Our last task is to compute the same angle $\tilde{\varphi}$ for the second photon at its arrival time $t_{\tilde{0}}=t_{0}+T_{D}$:
$\cos\tilde{\varphi}=\,\frac{\tilde{\vec{x}}^{\prime}_{\perp\tilde{0}}\cdot\vec{u}_{x\tilde{0}}}{\sqrt{\tilde{\vec{x}}^{\prime}_{\perp\tilde{0}}\cdot\tilde{\vec{x}}^{\prime}_{\perp\tilde{0}}}}\,=\,\frac{\tilde{A}}{\sqrt{\tilde{A}^{2}+\tilde{B}^{2}}}\,\sim\,\frac{\,\frac{A}{N}\,+\alpha}{\sqrt{\,\frac{A^{2}+B^{2}}{N^{2}}\,-2\,\frac{C}{N}\,\gamma}}\,.$ (49)
At this point we must resist a further linearization in $\alpha$ and $\gamma$. We will understand this necessity in the next section.
5 Linearizing the scale factors
We known that the anisotropy of the universe cannot be large and we will concentrate on axial Bianchi I universes close to an underlying maximally symmetric one with only one scale factor $a_{F}(t)$, $\cdot_{F}$ for Friedman. It is convenient to parameterize the small deviation $\eta(t)$ in the following way:
$a=\vcentcolon a_{F}\,[1-{\textstyle\frac{1}{2}}\eta],\quad c=\vcentcolon a_{F}\,[1+\eta],\quad\quad\quad\eta_{0}=0,\quad H\sim H_{F}-{\textstyle\frac{1}{2}}\eta^{\prime},\quad H_{c}\sim H_{F}+\eta^{\prime},$ (50)
where we indicate the linearization in $\eta$ and $V$ (neglecting cross-terms $\eta V$) also by $\sim$. Then we have
$W$ $\sim$ $\,\frac{a_{F}}{N}\left[1-{\textstyle\frac{1}{2}}\left(1-3\cos^{2}\theta\right)\eta\right],$ (51)
$\underline{z}_{0}+1$ $\sim$ $\,\frac{a_{F0}}{a_{Fe}}\left[1+{\textstyle\frac{1}{2}}\left(1-3\cos^{2}\theta\right)\eta_{e}+\cos\varphi\,\sin\theta\,V_{x}+\cos\theta\,V_{z}\right],$ (52)
[…]
$I$ $\sim$ $\chi_{e0}+O(\eta),\quad\quad J\ \sim\ \chi_{e0}+O(\eta),$ (54)
$\alpha$ $\sim$ $\frac{T_{D}}{a_{F0}\chi_{e0}}\,(3\,\cos\varphi\,\sin\theta\,\cos^{2}\theta\,\eta_{e}+[1-\cos^{2}\varphi\,\sin^{2}\theta]\,V_{x}$ (55)
$\hskip 42.67912pt-\cos\varphi\,\sin\theta\,\cos\theta\,V_{z}\,),$
$\gamma$ $\sim$ $-\sin\theta\,\frac{T_{D}}{a_{F0}\chi_{e0}}\left(3\,\sin\theta\,\cos\theta\,\eta_{e}+\cos\varphi\,\cos\theta\,V_{x}-\sin\theta\,V_{z}\right),$ (56)
where $\chi_{e0}$ is the conformal time of flight of the photon = the comoving geodesic ‘distance’ between $\vec{x}_{e}$ and $\vec{x}_{0}$ in the underlying Friedman universe,
$\chi_{e0}\vcentcolon=\int_{t_{e}}^{t_{0}}\,\frac{1}{a_{F}}\,\quad\text{and}\quad\vec{V}=a_{F0}\vec{v}.$ (57)
Consider first the drift with respect to $\varphi$, equation (49), and linearize it in $\eta$ and $V$, but not in $T_{D}/\!\!\left\{a_{F0}\chi_{e0}\right\}$:
$\cos\tilde{\varphi}\sim\hskip 409.71968pt$ (58)
$\frac{\cos\varphi+[3\cos\varphi\,\cos^{2}\!\theta\,\eta_{e}+(1-\cos^{2}\!\varphi\,\sin^{2}\!\theta)\!\!\left\{V_{x}/\sin\theta\right\}-\cos\varphi\,\cos\theta\,V_{z}]\,T_{D}/\!\!\left\{a_{F0}\chi_{e0}\right\}}{\sqrt{1+2\cos\theta[3\cos\theta\,\eta_{e}+\cos\varphi\,\cos\theta\left\{V_{x}/\sin\theta\right\}-V_{z}]\,T_{D}/\!\!\left\{a_{F0}\chi_{e0}\right\}}}\,.$
For $\sin\theta=0$, the angle $\varphi$ is ill-defined and equation (58) is singular:
[…]
If not only $V\ll 1$, but also |$|\{V_{x}/\sin\theta\}|\ll 1$, we may further [9] linearize $\cos\tilde{\varphi}$:
$\cos\tilde{\varphi}$ $\sim$ $\cos\varphi\left[1+\,\frac{\sin^{2}\varphi}{\cos\varphi}\,\frac{V_{x}}{\sin\theta}\,\frac{T_{D}}{a_{F0}\chi_{e0}}\right].$ (60)
In this case the drift with respect to the angle $\varphi$ is,
[…]
$\delta\varphi$ $\sim$ $\frac{-1}{\sin\varphi}\,\,\delta\cos\varphi\sim-\,{\sin\varphi}\,\frac{V_{x}}{\sin\theta}\,\frac{T_{D}}{a_{F0}\chi_{e0}}\,.$ (62)
In particular if $V_{x}$ vanishes, then $\varphi$ is constant and the drift remains in the plane defined by the $z$-axis and this constant angle $\varphi$.
The definition of the angle $\theta$ does not involve the projection into the $xy$-plane and linearizing the $\theta$-drift is straight-forward:
$\cos\tilde{\theta}$ $\sim$ $\cos\theta\left[1-\sin^{2}\theta\left(3\left({\textstyle\frac{1}{2}}a_{F0}\chi_{e0}\,\eta^{\prime}_{0}+\eta_{e}\right)+\cos\varphi\,\frac{V_{x}}{\sin\theta}\,-\,\frac{V_{z}}{\cos\theta}\,\right)\frac{T_{D}}{a_{F0}\chi_{e0}}\right],$ (63)
$\delta\cos\theta$ $\sim$ $-\sin^{2}\theta\,\cos\theta\left(3\left({\textstyle\frac{1}{2}}a_{F0}\chi_{e0}\,\eta^{\prime}_{0}+\eta_{e}\right)+\cos\varphi\,\frac{V_{x}}{\sin\theta}\,-\,\frac{V_{z}}{\cos\theta}\right)\frac{T_{D}}{a_{F0}\chi_{e0}},$ (64)
$\delta\theta$ $\sim$ $\left(3\sin\theta\,\cos\theta\left({\textstyle\frac{1}{2}}a_{F0}\chi_{e0}\,\eta^{\prime}_{0}+\eta_{e}\right)+\cos\varphi\,\cos\theta\,V_{x}-\sin\theta\,V_{z}\right)\,\frac{T_{D}}{a_{F0}\chi_{e0}}\,.$ (65)
Note that the first term, $\eta^{\prime}_{0}$, in the drift (65) does not depend on the redshift of the source. For a dominating expansion in the $z$-direction today, $H_{c0}-H_{0}\sim{\textstyle\frac{3}{2}}\eta^{\prime}_{0}>0$, this first term produces a drift away from the $z$-axis.
The second term, $\eta_{e}$, in the drift (65) produces the opposite effect. Indeed, for a dominating expansion in the $z$-direction from emission until today, $\eta^{\prime}$ is positive and, since we have set $\eta_{0}$ to zero, $\eta_{e}$ is negative.
If we adopt Einstein’s equations with cosmological constant and dust, then $a_{F}$ and $\eta$ are given by,
$a_{F}(t)=a_{F0}\left(\frac{\cosh\left(\sqrt{3\Lambda}\,t\right)-1}{\cosh\left(\sqrt{3\Lambda}\,t_{0}\right)-1}\right)^{1/3},\hskip 11.38109ptt_{0}={\textstyle\frac{2}{3}}\,\frac{\text{arc}\tanh\,\sqrt{\Omega_{\Lambda 0}}}{\sqrt{\Omega_{\Lambda 0}}}\,\frac{1}{H_{F0}}\,,\hskip 11.38109pt\Omega_{\Lambda 0}\vcentcolon=\,\frac{\Lambda}{3\,H_{F0}^{2}}\,,$ (66)
and
$\eta^{\prime}=\eta^{\prime}_{0}\,\frac{a_{F0}^{3}}{a_{F}^{3}}\quad\text{implying}\quad\eta(t)=-\,{\textstyle\frac{2}{3}}\,\frac{\eta^{\prime}_{0}}{H_{F0}}\,\frac{\sqrt{\Omega_{\Lambda 0}}}{1-\Omega_{\Lambda 0}}\left[\coth\left({\textstyle\frac{1}{2}}\sqrt{3\Lambda}\,t\right)-1/\sqrt{\Omega_{\Lambda 0}}\right].$ (67)
We take $a_{F0}=1$ Gyr, $\Omega_{\Lambda 0}=0.7$ and $H_{0}=72$ km/s/Mpc implying $13.581/H_{F0}=$ 13.58 Gyr, $t_{0}=13.09$ Gyr and $9.371/\sqrt{\Lambda}=$ 9.37 Gyr.
Let us compute the ratio $\eta_{e}/{(\textstyle\frac{1}{2}}a_{F0}\chi_{e0}\,\eta^{\prime}_{0})$ of the second term in the linearized drift equation (65) by the first term. For the redshift $\underline{z}_{0}=1$, we find $t_{e}=5.59$ Gyr and by numerical integration $\chi_{e0}=10.48$. Finally we obtain,
$\,\frac{\eta_{e}}{{\textstyle\frac{1}{2}}a_{F0}\chi_{e0}\,\eta^{\prime}_{0}}\,=-\,4.38.$ (68)
Our result disagrees with the result by Quercellini et al. [10][11] without peculiar velocity. They find – as the redshift ranges from zero to one – that the absolute value of the ratio is at most 7 %. They also find the opposite sign in front of the first term. The sign problem was noticed before by Marcori et al. [12], who also include peculiar velocities.
6 Conclusion
This second term in the drift equation (65), which is proportional to $\eta_{e}=\eta(t_{e})$ and which we find to be dominant, depends on the redshift of the source via the inverse of the direction-dependent relation (52). It is therefore phenomenologically more interesting for a fit of the Gaia catalogue of quasars than the first term. We look forward to this fit.
Acknowledgements: I thank André Tilquin and Galliano Valent for tens of years of joint research and fun. It is also a pleasure to thank Jeremy Darling for his constructive advice.